OpenStudy (anonymous):

0.0007 x 11,000/0.00024 Express your answer in the form of a scientific notation.

OpenStudy (hba):

Are u allowed to use a scientific calculator

OpenStudy (anonymous):

no, my teacher said and I dont have that calculator too

OpenStudy (anonymous):

@hba can u help me?

OpenStudy (anonymous):

@waterineyes can u help me please

OpenStudy (anonymous):

I am not able to post anything I don't know why..

OpenStudy (anonymous):

Draw tool is not working..

OpenStudy (anonymous):

can u send me through message

OpenStudy (anonymous):

or maybe just type in here @waterineyes

OpenStudy (anonymous):

\[\frac{7 \times 11000 \times 100000}{24 \times 10000} \implies \frac{77}{24}\times 10^4 \implies 3.2 \times 10^4\]

OpenStudy (anonymous):

Getting??

OpenStudy (anonymous):

my answer is 7.7x10 to the power of 0 /2.4 x10 to the power of 4

OpenStudy (anonymous):

There is one more method: 0.0007 can be written as : \(7 \times 10^{-4}\) 11000 can be written as : \(11 \times 10^3\) 0.00024 can be written as : \(24 \times 10^{-5}\) So, you will get: \[\frac{77 \times 10^{-1}}{24 \times 10^{-5}} \implies \frac{7.7 \times 10^0}{2.4 \times 10^{-4}}\] I got -4 in the power.. Is it right??

OpenStudy (anonymous):

i think

OpenStudy (anonymous):

Yes it is...

OpenStudy (anonymous):

thanks a lot @waterineyes

OpenStudy (anonymous):

Welcome dear..